Convert an absolute pressure of 6.64×105 N/m2 to gauge pressure in lb/in2.
Answer #1

absolute pressure =gauge pressure + atomspehric pressure

gauge pressure = absolute pressure - atomspehric pressure'

gauge pressure =  6.64×10^5 - 1.01 x 10^ 5 = 5.63 x 10^ 5 N/m^2

1 N/m^2 = 0.000145038 lb./ in ^2

5.63 x 10^ 5 = 81.656394 lb/in^2